# Phase separation of chemokinetic active particles

### v1

@@ A.1 Case of the BMR
$D_{\text{eff}}<\frac{\alpha n_{0}}{2D_{\mathrm{r}}}(\alpha n_{0}-\rho_{0}\zeta).$ (S11)
@@
$D_{\text{eff}}>\frac{\alpha n_{0}}{2D_{\mathrm{r}}}(\alpha n_{0}-\rho_{0}\zeta)$ (S12)
@@
$D_{\text{eff}}<-\frac{\kappa\lambda\rho_{0}}{D_{\text{c}}}$ (S13)
must be satisfied simultaneously, which is impossible. Thus, only Eq. (S11) determines the S phase, which can also be written as
$D<\frac{\rho_{0}\zeta(\alpha n_{0}-\rho_{0}\zeta)}{D_{\mathrm{r}}}.$ (S14)
@@
Meanwhile, in the absence of particle-chemical coupling ($\lambda=0$), $h(q^{2})<0$ holds for some range of $q^{2}$ if and only if $h^{\prime}(0)=D_{\text{eff}}D_{\text{c}}<0$. This implies that the vanilla model exhibits MIPS for $D_{\text{eff}}<0$ (which is clearly more stringent than Eq. (S11) for the BMR), from which Eq. (12) is derived.
A.2 Case of the AMR
@@
Let us first determine the boundaries of the H phase, which requires that the real part of $\omega_{\pm}$ is negative for all $q^{2}$. Due to the quadratic nature of Eq. (S8), the condition $\mathrm{Re}~{}\omega_{\pm}<0$ is fulfilled when both $\mathcal{A}\mathcal{D}-\mathcal{B}\mathcal{C}>0$ and $\mathcal{A}+\mathcal{D}<0$ are true.
@@
$h(q^{2})$ $=\kappa D_{\mathrm{c}}q^{4}+[D_{\mathrm{eff}}D_{\mathrm{c}}+\kappa\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)]q^{2}$
$\quad+\lambda\rho_{0}D_{\mathrm{eff}}(2\alpha n_{0}-\rho_{0}\zeta)$
$\quad-\frac{\lambda\rho_{0}}{2D_{\mathrm{r}}}\alpha n_{0}(\alpha n_{0}-\rho_{0}\zeta)(\alpha n_{0}-2\rho_{0}\zeta)$ (S15)
@@
$D_{\mathrm{eff}}>\max$ $\bigg{[}-\frac{\kappa\lambda\rho_{0}}{D_{\mathrm{c}}}(2\alpha n_{0}-\rho_{0}\zeta),$
$\quad\frac{\alpha n_{0}(\alpha n_{0}-\rho_{0}\zeta)(\alpha n_{0}-2\rho_{0}\zeta)}{2D_{\mathrm{r}}(2\alpha n_{0}-\rho_{0}\zeta)}\bigg{]}\;.$ (S16)
@@
$-\frac{1}{4\kappa D_{\mathrm{c}}}\left[D_{\mathrm{eff}}D_{\mathrm{c}}+\kappa\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)\right]^{2}>0\;,$ (S17)
@@
$\kappa q^{4}+(D_{\text{eff}}+D_{\mathrm{c}})q^{2}+\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)>0\;.$ (S18)
@@
$\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)-\frac{(D_{\text{eff}}+D_{\mathrm{c}})^{2}}{4\kappa}>0\;.$ (S19)
@@
$D_{\text{eff}}>-D_{\text{c}}-2\sqrt{\kappa\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)}\;.$ (S20)
Based on the above, one can fully determine the boundaries of the H phase. Here we refrain from giving the explicit expressions for the corresponding values of $D_{\mathrm{eff}}$, which is a rather tedious task. Instead, we note that setting $\lambda>0$ broadens the ranges of $D_{\mathrm{eff}}$ associated with the H phase. This can be seen through the following argument. First, we observe that Eqs. (A.2) and (A.2) imply
$D_{\text{eff}}>\frac{\alpha n_{0}(\alpha n_{0}-\rho_{0}\zeta)(\alpha n_{0}-2\rho_{0}\zeta)}{2D_{\mathrm{r}}(2\alpha n_{0}-\rho_{0}\zeta)}\;.$ (S21)
Using this inequality in Eq. (A.2), we obtain
$h(q^{2})\geq\kappa D_{\mathrm{c}}q^{4}+D_{\mathrm{eff}}D_{\mathrm{c}}q^{2}\;,$ (S22)
with the equality holding if and only if $\lambda=0$. This means that $\mathcal{A}\mathcal{D}-\mathcal{B}\mathcal{C}>0$ becomes weaker for $\lambda>0$ than for $\lambda=0$. Similarly, Eq. (S20) shows that $\mathcal{A}+\mathcal{D}<0$ becomes weaker for $\lambda>0$ than for $\lambda=0$. Since both inequalities are weakened by positive $\lambda$, we conclude that the AMR suppresses the MIPS and widens the H phase, just as expected in Sec. II.
@@
Table 1: Ordering of $q_{\text{l,r}}^{2}$ and $q_{\pm}^{2}$. $q_{\pm}^{2}$ are bold-faced to emphasize the changes when transitioning from $y<4x(1-2x)$ to $y>4x(1-2x)$.
@@
$q_{c}^{2}$ $\equiv-\frac{D_{\text{eff}}-D_{\mathrm{c}}}{2\kappa}$ (S23)
$\,\,\,\,\,\,+\frac{\sqrt{(D_{\text{eff}}-D_{\mathrm{c}})^{2}+4\kappa\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)}}{2\kappa}\;.$
@@
Figure S1: Illustrations of intervals satisfying $f(q^{2})<g(q^{2})$. The black solid line represents the quadratic function $f(q^{2})$, and the blue dashed line corresponds to the linear function $g(q^{2})$. The points of tangency at $q_{\mathrm{l}}^{2}$, $q_{\mathrm{r}}^{2}$, and $q_{\mathrm{c}}^{2}$ are marked by the circle, the triangle, and the square, respectively. The intervals of $q^{2}$ with $f(q^{2})<g(q^{2})$ are indicated by thick red horizontal segments.
@@
$3\kappa q^{4}+(D_{\text{eff}}-D_{\text{c}})q^{2}+\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)=0\;.$ (S25)
@@
$(D_{\text{eff}}-D_{\text{c}})^{2}-12\kappa\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)\geq 0\;.$ (S26)
@@
$q_{\text{l},{\text{r}}}^{2}$ $=-\frac{D_{\text{eff}}-D_{\mathrm{c}}}{6\kappa}$ (S27)
±$\,\,\,\,\,\,\pm\frac{\sqrt{(D_{\text{eff}}-D_{\mathrm{c}})^{2}-12\kappa\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)}}{6\kappa}\;,$
where the plus (minus) sign corresponds to $q_{\mathrm{r}}$ ($q_{\mathrm{l}})$. We note that both $q_{\mathrm{l}}^{2}$ and $q_{\mathrm{r}}^{2}$ are positive if they have real values. As illustrated in Fig. S1, depending on the slope of $g(q^{2})$ (indicated by the dashed line), the intervals of $q^{2}$ satisfying the condition $f(q^{2})<g(q^{2})$ (indicated by the thick horizontal lines) vary. When the slope of $g(q^{2})$ is less than that of the tangent line crossing $(q_{\mathrm{l}}^{2},f(q_{\mathrm{l}}^{2}))$ (indicated by the circle), there exists only a single interval satisfying $f(q^{2})<g(q^{2})$. If the slope of $g(q^{2})$ is greater than that of the tangent line crossing $(q_{\mathrm{r}}^{2},f(q_{\mathrm{r}}^{2}))$, the inequality is satisfied everywhere. In the intermediate regime, there are two separate intervals satisfying the inequality, one around $q_{\mathrm{l}}^{2}$ and the other around $q_{\mathrm{c}}^{2}$. As discussed earlier, $\omega_{\pm}$ have nonzero imaginary parts in these intervals.
@@
$q_{\pm}^{2}$ $=-\frac{D_{\text{eff}}+D_{\mathrm{c}}}{2\kappa}$ (S28)
±$\quad\pm\frac{\sqrt{(D_{\text{eff}}+D_{\mathrm{c}})^{2}-4\kappa\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)}}{2\kappa}\;.$
@@
This equation is useful for checking when the ordering of $q_{\text{l,r}}^{2}$ and $q_{\pm}^{2}$ changes. See Table 1 for all possible orderings. For most cases, we have $q_{\mathrm{l}}^{2}<q_{-}^{2}<q_{+}^{2}$. Then the interval $(q_{-}^{2},q_{+}^{2})$ overlaps with nonzero imaginary parts of $\omega_{\pm}$ if and only if $\Delta(q^{2})\equiv f(q^{2})-g(q^{2})$ becomes negative at $q^{2}=q_{-}^{2}$ or $q_{+}^{2}$. In contrast, for $x<1/5$ and $y>4x(1-2x)$, the ordering changes to $q_{-}^{2}<q_{\mathrm{l}}^{2}<q_{+}^{2}$. It is then possible that $(q_{-}^{2},q_{+}^{2})$ fully contains an interval around $q_{\mathrm{l}}^{2}$ with nonzero imaginary parts of $\omega_{\pm}$, e.g., the left one of the two intervals indicated by thick horizontal segments in the middle panel of Fig. S1. For this particular case, the O phase requires that at least one among $\Delta(q_{-}^{2})$, $\Delta(q_{\mathrm{l}}^{2})$, and $\Delta(q_{+}^{2})$ must be negative.
@@
Finally, if Eq. (S26) is not satisfied, then $f(q^{2})<g(q^{2})$ can hold only across a single continuous interval. Then, the O phase only requires that $\Delta(q_{+}^{2})$ or $\Delta(q_{-}^{2})$ be negative. By combining all the conditions discussed so far, the boundaries of the O phase can be fully determined.

### v2

@@ B.0.1 Case of the BMR
$D_{\text{eff}}<\frac{\alpha n_{0}}{2D_{\mathrm{r}}}(\alpha n_{0}-\rho_{0}\zeta).$ (A29)
@@
$D_{\text{eff}}>\frac{\alpha n_{0}}{2D_{\mathrm{r}}}(\alpha n_{0}-\rho_{0}\zeta)$ (A30)
@@
$D_{\text{eff}}<-\frac{\kappa\lambda\rho_{0}}{D_{\text{c}}}$ (A31)
must be satisfied simultaneously, which is impossible. Thus, only Eq. (A29) determines the S phase, which can also be written as
$D<\frac{\rho_{0}\zeta(\alpha n_{0}-\rho_{0}\zeta)}{D_{\mathrm{r}}}.$ (A32)
@@
Meanwhile, in the absence of particle–chemical coupling ($\lambda=0$), $h(q^{2})<0$ holds for some range of $q^{2}$ if and only if $h^{\prime}(0)=D_{\text{eff}}D_{\text{c}}<0$. This implies that the vanilla model exhibits MIPS for $D_{\text{eff}}<0$ (which is clearly more stringent than Eq. (9) in the main text for the BMR), from which Eq. (11) in the main text is derived.
B.0.2 Case of the AMR
@@
Let us first determine the boundaries of the H phase, which requires that the real part of $\omega_{\pm}$ is negative for all $q^{2}$. Due to the quadratic nature of Eq. (A26), the condition $\mathrm{Re}~{}\omega_{\pm}<0$ is fulfilled when both $\mathcal{A}\mathcal{D}-\mathcal{B}\mathcal{C}>0$ and $\mathcal{A}+\mathcal{D}<0$ are true.
@@
$h(q^{2})$ $=\kappa D_{\mathrm{c}}q^{4}+[D_{\mathrm{eff}}D_{\mathrm{c}}+\kappa\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)]q^{2}+\lambda\rho_{0}D_{\mathrm{eff}}(2\alpha n_{0}-\rho_{0}\zeta)$
$\quad-\frac{\lambda\rho_{0}}{2D_{\mathrm{r}}}\alpha n_{0}(\alpha n_{0}-\rho_{0}\zeta)(\alpha n_{0}-2\rho_{0}\zeta)$ (A33)
@@
$D_{\mathrm{eff}}>\max\bigg{[}-\frac{\kappa\lambda\rho_{0}}{D_{\mathrm{c}}}(2\alpha n_{0}-\rho_{0}\zeta),\frac{\alpha n_{0}(\alpha n_{0}-\rho_{0}\zeta)(\alpha n_{0}-2\rho_{0}\zeta)}{2D_{\mathrm{r}}(2\alpha n_{0}-\rho_{0}\zeta)}\bigg{]}\;.$ (A34)
@@
$-\frac{1}{4\kappa D_{\mathrm{c}}}\left[D_{\mathrm{eff}}D_{\mathrm{c}}+\kappa\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)\right]^{2}>0\;,$ (A35)
@@
$\kappa q^{4}+(D_{\text{eff}}+D_{\mathrm{c}})q^{2}+\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)>0\;.$ (A36)
@@
$\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)-\frac{(D_{\text{eff}}+D_{\mathrm{c}})^{2}}{4\kappa}>0\;.$ (A37)
@@
$D_{\text{eff}}>-D_{\text{c}}-2\sqrt{\kappa\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)}\;.$ (A38)
Based on the above, one can fully determine the boundaries of the H phase. Here we refrain from giving the explicit expressions for the corresponding values of $D_{\mathrm{eff}}$, which is a rather tedious task. Instead, we note that setting $\lambda>0$ broadens the ranges of $D_{\mathrm{eff}}$ associated with the H phase. This can be seen through the following argument. First, we observe that Eqs. (A34) and (B.0.2) imply
$D_{\text{eff}}>\frac{\alpha n_{0}(\alpha n_{0}-\rho_{0}\zeta)(\alpha n_{0}-2\rho_{0}\zeta)}{2D_{\mathrm{r}}(2\alpha n_{0}-\rho_{0}\zeta)}\;.$ (A39)
Using this inequality in Eq. (B.0.2), we obtain
$h(q^{2})\geq\kappa D_{\mathrm{c}}q^{4}+D_{\mathrm{eff}}D_{\mathrm{c}}q^{2}\;,$ (A40)
with the equality holding if and only if $\lambda=0$. This means that $\mathcal{A}\mathcal{D}-\mathcal{B}\mathcal{C}>0$ becomes weaker for $\lambda>0$ than for $\lambda=0$. Similarly, Eq. (A38) shows that $\mathcal{A}+\mathcal{D}<0$ becomes weaker for $\lambda>0$ than for $\lambda=0$. Since both inequalities are weakened by positive $\lambda$, we conclude that the AMR suppresses the MIPS and widens the H phase.
@@
Table A1: Ordering of $q_{\text{l,r}}^{2}$ and $q_{\pm}^{2}$. $q_{\pm}^{2}$ are bold-faced to emphasize the changes when transitioning from $y<4x(1-2x)$ to $y>4x(1-2x)$.
@@
$q_{c}^{2}\equiv-\frac{D_{\text{eff}}-D_{\mathrm{c}}}{2\kappa}+\frac{\sqrt{(D_{\text{eff}}-D_{\mathrm{c}})^{2}+4\kappa\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)}}{2\kappa}\;.$ (A41)
@@
Figure A1: Wave number intervals leading to oscillating fluctuations. The black solid line represents the quadratic function $f(q^{2})$, and the blue dashed line corresponds to the linear function $g(q^{2})$. The points of tangency at $q_{\mathrm{l}}^{2}$, $q_{\mathrm{r}}^{2}$, and $q_{\mathrm{c}}^{2}$ are marked by the circle, the triangle, and the square, respectively. The intervals of $q^{2}$ with $f(q^{2})<g(q^{2})$, which produce oscillating fluctuations, are indicated by thick red horizontal segments.
@@
$3\kappa q^{4}+(D_{\text{eff}}-D_{\text{c}})q^{2}+\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)=0\;.$ (A43)
@@
$(D_{\text{eff}}-D_{\text{c}})^{2}-12\kappa\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)\geq 0\;.$ (A44)
@@
$q_{\text{l},{\text{r}}}^{2}=-\frac{D_{\text{eff}}-D_{\mathrm{c}}}{6\kappa}\pm\frac{\sqrt{(D_{\text{eff}}-D_{\mathrm{c}})^{2}-12\kappa\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)}}{6\kappa}\;,$ (A45)
where the plus (minus) sign corresponds to $q_{\mathrm{r}}$ ($q_{\mathrm{l}})$. We note that both $q_{\mathrm{l}}^{2}$ and $q_{\mathrm{r}}^{2}$ are positive if they have real values. As illustrated in Fig. A1, depending on the slope of $g(q^{2})$ (indicated by the dashed line), the intervals of $q^{2}$ satisfying the condition $f(q^{2})<g(q^{2})$ (indicated by the thick horizontal lines) vary. When the slope of $g(q^{2})$ is less than that of the tangent line crossing $(q_{\mathrm{l}}^{2},f(q_{\mathrm{l}}^{2}))$ (indicated by the circle), there exists only a single interval satisfying $f(q^{2})<g(q^{2})$. If the slope of $g(q^{2})$ is greater than that of the tangent line crossing $(q_{\mathrm{r}}^{2},f(q_{\mathrm{r}}^{2}))$, the inequality is satisfied everywhere. In the intermediate regime, there are two separate intervals satisfying the inequality, one around $q_{\mathrm{l}}^{2}$ and the other around $q_{\mathrm{c}}^{2}$. As discussed earlier, $\omega_{\pm}$ have nonzero imaginary parts in these intervals.
@@
$q_{\pm}^{2}=-\frac{D_{\text{eff}}+D_{\mathrm{c}}}{2\kappa}\pm\frac{\sqrt{(D_{\text{eff}}+D_{\mathrm{c}})^{2}-4\kappa\lambda\rho_{0}(2\alpha n_{0}-\rho_{0}\zeta)}}{2\kappa}\;.$ (A46)
@@
This equation is useful for checking when the ordering of $q_{\text{l,r}}^{2}$ and $q_{\pm}^{2}$ changes. See Table A1 for all possible orderings. For most cases, we have $q_{\mathrm{l}}^{2}<q_{-}^{2}<q_{+}^{2}$. Then the interval $(q_{-}^{2},q_{+}^{2})$ overlaps with nonzero imaginary parts of $\omega_{\pm}$ if and only if $\Delta(q^{2})\equiv f(q^{2})-g(q^{2})$ becomes negative at $q^{2}=q_{-}^{2}$ or $q_{+}^{2}$. In contrast, for $x<1/5$ and $y>4x(1-2x)$, the ordering changes to $q_{-}^{2}<q_{\mathrm{l}}^{2}<q_{+}^{2}$. It is then possible that $(q_{-}^{2},q_{+}^{2})$ fully contains an interval around $q_{\mathrm{l}}^{2}$ with nonzero imaginary parts of $\omega_{\pm}$, e.g., the left one of the two intervals indicated by thick horizontal segments in the middle panel of Fig. A1. For this particular case, the O phase requires that at least one among $\Delta(q_{-}^{2})$, $\Delta(q_{\mathrm{l}}^{2})$, and $\Delta(q_{+}^{2})$ must be negative.
@@
Finally, if Eq. (A44) is not satisfied, then $f(q^{2})<g(q^{2})$ can hold only across a single continuous interval. Then, the O phase only requires that $\Delta(q_{+}^{2})$ or $\Delta(q_{-}^{2})$ be negative. By combining all the conditions discussed so far, the boundaries of the O phase can be fully determined.
Appendix C Effects of upper bound on self-propulsion
In this note, we examine the case where the self-propulsion of the chemokinetic ABPs in the AMR saturates to a finite upper bound as the chemical concentration $n$ is increased. This is imposed by
$\tilde{v}(n)=\alpha n_{\mathrm{b}}\tanh\left(\frac{n}{n_{\mathrm{b}}}\right),$ (A48)
where $n_{\mathrm{b}}$ indicates the concentration scale at which $\tilde{v}(n)$ saturates, in place of $\tilde{v}(n)=\alpha n$. The same assumption was used in the literature [64] to account for the maximum self-propulsion achievable even when the chemical is abundant. In Fig. A2, we show how the giant cluster formed by the ABPs change as a function of $n_{\mathrm{b}}$. When $n_{\mathrm{b}}$ is reduced, the giant cluster becomes larger, as the self-propulsion increases within the clusters to a much less degree. In contrast, as $n_{\mathrm{b}}$ is increased, the upper limit on the self-propulsion is relaxed, making $\tilde{v}(n)\simeq\alpha n$ a better approximation. Thus, the monotonic decrease of the giant cluster fraction with increasing $n_{\mathrm{b}}$ confirms our view that the chemokinetic speedup inside the cluster is responsible for the supression of MIPS in the AMR.
Figure A2: Fraction of the giant cluster formed by the active Brownian particles with repulsive interactions in the presence of an upper bound on self-propulsion. The saturation of the self-propulsion speed is given by Eq. (A48). The circles and the error bars indicate the average and the standard deviation, respectively, obtained using six different simulations. All parameters, except for $n_{\mathrm{b}}$, are identical to those used in Fig. 2 in the main text.
Figure A3: Evolution of particle density $\rho(\mathbf{r},t)$ and chemical concentration $n(\mathbf{r},t)$ in the Basal Metabolic Regime of the hydrodynamic model. We use (a) $n_{0}=18$, $\zeta=6$ and (b) $n_{0}=6$, $\zeta=6$, while the other parameters are identical to those used in Fig. 3(a) in the main text. Even as the global particle density remains the same ($\rho_{0}=1/2$), the system can be made to form either (a) clusters or (b) bubbles by changing the global chemical concentration $n_{0}$. See Supplementary Movies 5A and 5B for animated versions of (a) and (b), respectively.
Appendix D Evolution of patterns in hydrodynamic simulations
In this note, we present demonstrations of how the particle density and the chemical concentration of the hydrodynamic model evolve to form clusters or bubbles. We first show how the phase-separated patterns in the BMR evolve with time, see Fig. A3. For the vanilla MIPS, whether phase separation occurs via bubble or cluster formation depends on the global density $\rho_{0}$ of particles. However, by introducing chemokinetic effects, the particles can be made to form clusters [as shown in Fig. A3(a)] or bubbles [as shown in Fig. A3(b)] even at the same value of $\rho_{0}$ by adjusting the global chemical concentration $n_{0}$. This is thanks to the strong anticorrelation between $\rho(\mathbf{r},t)$ and $n(\mathbf{r},t)$, which makes the chemicals form clusters (bubbles) where the particles form bubbles (clusters). Since large (small) $n_{0}$ favors bubble (cluster) formation of the chemicals, the particles are driven to form clusters (bubbles). We also note that the length scale of the bubbles and clusters gradually increases in time, which suggests the coarsening dynamics driven by the Ostwald process.
Figure A4: Evolution of particle density $\rho(\mathbf{r},t)$ and chemical concentration $n(\mathbf{r},t)$ in the hydrodynamic simulation of the Active Metabolic Regime, starting from the initial state with two particle clusters and the uniform concentration. The parameters are given by $n_{0}=20$, $\zeta=25$, $D_{\mathrm{c}}=0.1$, and $\lambda=0.5$, which corresponds to the S phase. See Supplementary Movie 6 for an animated version.
Next, we turn to the AMR and examine how the microphase separation occurs in the S phase, starting from a configuration of two large particle clusters and the uniform chemical distribution. As shown in Fig. A4, these large particle clusters quickly disintegrate into smaller clusters (including some ring-like structures), while the chemical develops a pattern that overlaps almost exactly with the particle distribution. Then these intermediate particle–chemical clusters gradually shrink, letting the other particle–chemical clusters grow at their expense until all clusters have roughly the same size. This suggests the reverse Ostwald process underlying the microphase separation, which was also found to play a crucial role in the microphase separation of generic single-species scalar active matter [16].
Appendix E Extra results for quorum-sensing active particles